Problem Solving:

1.     What is the electric field strength at a point that is 0.25 m to the left of a small sphere with a charge of -3.8 x 10 -6 C?

 

 

 

2.     A +2.0 mC charge is placed in an electric field at a point with a magnitude electric field equivalent to 4.23 x 105N/C. Determine the magnitude of the electrostatic force acting on the charge.

 

 

 

3.     A sphere has a 400 N/C electric field. A small sphere is 0.20 away from the sphere. What is the charge in the sphere?

 

1
Expert's answer
2021-04-28T20:18:55-0400

1. The magnitude of the electric field is


"E=\frac{1}{4\pi\epsilon_0}\frac{q}{r^2}=546.4·10^3\text{ N/C}."

2. The magnitude of the electrostatic force acting on the charge is


"F=Eq=(4.23·10^5)(2·10^{-3})=846\text{ N}."

3. If 0.20 m is the distance at which the field of 400 N/C was measured, and this distance is greater than the radius of the sphere, the charge can be found from the equation for the field (see problem (1)):


"q=4\pi\epsilon_0r^2E=1.78·10^{-9}\text{ C}."